# SIR: Multi-view Inverse Rendering with Decomposable Shadow for Indoor Scenes

### 25 Feb 2024 v2

@@ B.1 Implemetaions of SIR
Phase 2: Hard shadow and irradiance field. For the representation of irradiance $I_{ir}$ and hard shadow $S_{hard}$, we use two separate MLPs, each consisting of 4 layers of 256 hidden units with a rectified linear unit (ReLU) activation function. In addition, we encode the input surface position $\hat{\textbf{x}}$ with 10 levels of periodic functions, respectively, before feeding them into our network. We use the incoming diffuse lighting $L_{i,d}$ in Sec. 3.2 and hard shadow $\xi$ in Sec. 3.3 as the pseudo-GT to supervise the predicted irradiance $I_{ir}$ and hard shadow $S_{hard}$ via $\ell_{1}$ loss. We use 512 sample directions over hemisphere $\Omega^{+}$ to precompute the diffuse lighting $L_{i,d}$ and hard shadow $\xi$ at sampled surface points. We set the threshold $\mu$ as 4 in synthetic scenes and 3 in real-world scenes to obtain the hard shadow $\xi$ by distinguishing HDR radiance intensity between the main light and environment light. We optimize our irradiance and hard shadow MLPs for 10k iterations with a batch size of 256 in this stage, which takes about 7 hours for a scene.
Phase 3: BRDF and soft shadow field. In the section of material estimation (Sec. 3.4), the BRDF MLPs include four fully connected layers, each with 512 hidden units and ReLU activation functions. Following these four layers, the BRDF feature network is divided into separate albedo and roughness layers with 512 hidden units and Sigmoid activation. Specifically, albedo layer outputs the albedo term $\hat{A}\in\mathbb{R}^{3}$ and roughness layer outputs the roughness term $\hat{R}\in\mathbb{R}^{1}$ at surface position $\hat{\textbf{x}}$. We use 64 sample directions (8 $\times$ 16) over hemisphere $\Omega^{+}$ at sampled surface points for the specular environment map, and then when we use the specular environment map as specular incoming lighting with specular BRDF term $f_{s}$ to compute the rendering equation. We use an important sampling method to sample the incoming light radiance by interpolating the specular environment map based on nvdiffrast [56], the important sampling number is 128.
@@ B.2 Implemetaions of Baselines
Appendix C Details of Experiments
C.1 Additional Results on Synthetic Dataset
We present additional qualitative (see Fig. 6 and 7) and quantitative (see Table 5,  6,  7,  8,  9,  10, and  11) comparison results on the synthetic dataset including 6 scenes. Compared to other inverse rendering methods, our approach excels in estimating albedo. Crucially, it effectively separates shadows and ensures that they are not incorrectly included in the albedo, especially under indoor lighting conditions. Note that PhySG [32] and InvRender [33] directly use our trained SDF-based radiance field as their geometry backbone. Since the predicted roughness of PhySG is global, it cannot converge on our indoor scene dataset. Therefore, we ignore the comparison of the roughness MSE metric with PhySG.
C.2 Additional Results on Real Dataset
Our comparative analysis on the real-world dataset, constrained to view synthesis metrics due to the unavailability of real scene material ground truths, is presented in Table LABEL:tab:exp_realworld_res. Although our method is slightly inferior to IBL-NeRF in terms of PSNR, it significantly outperforms other methods regarding SSIM and LPIPS metrics. Fig. 8 provides additional qualitative results that highlight our exceptional decomposition ability. We effectively disentangle albedo and shadow components, achieving physical plausibility in our results.
C.3 Additional Results for Ablation studies
We showcase the effectiveness of our material optimization on synthetic datasets in Table 12131415161718, and 19. As described in Sec. 4.4 in the main paper, Without shadow terms significantly impacts albedo estimation, leading to incorrect albedo outputs that erroneously incorporate baked shadows; Without the soft shadow term, residual shadows remain baked into the albedo, and due to the binary nature of hard shadows, they fail to accurately represent physically plausible shadow elements; Without the incorporation of an instance-level albedo regularizer, there would likely be a lack of illumination consistency in the albedo for individual instances, potentially leading to less accurate synthesis results. We show more qualitative results in Fig. 9 and 10.
@@ C.3 Additional Results for Ablation studies
Figure 7: Qualitative results of all methods on the synthetic dataset (Study Room).
@@
Figure 10: Qualitative results of ablation study on the synthetic dataset (left: Reception, right: Study Room).
[…]
Dinning Office Avg. Dinning Office Avg. Dinning Office Avg.
IBL-NeRF [36] 26.2340 22.3221 24.2781 0.7968 0.7013 0.7491 0.4285 0.6165 0.5225

### 09 Apr 2024 v3

@@ 0.B.1 Implemetaions of SIR
Phase 2: Hard shadow and irradiance field. For the representation of irradiance $I$ and hard shadow $S_{hard}$, we use two separate MLPs, each consisting of 4 layers of 256 hidden units with a rectified linear unit (ReLU) activation function. In addition, we encode the input surface position $\hat{\textbf{x}}$ with 10 levels of periodic functions, respectively, before feeding them into our network. We use the incoming diffuse lighting $L_{i,d}$ in Sec. 3.3 and hard shadow $\xi$ in Sec. 3.4 as the pseudo ground truths to supervise the predicted irradiance $I$ and hard shadow $S_{hard}$ via $\ell_{1}$ loss. We use 512 sample directions over hemisphere $\Omega^{+}$ to precompute the diffuse lighting $L_{i,d}$ and hard shadow $\xi$ at sampled surface points. To reduce the number of samples, we set the threshold $\mu$ as 4 in synthetic scenes and 3 in real-world scenes to obtain the hard shadow $\xi$ by distinguishing HDR radiance intensity between the main light and environment light. We optimize irradiance and hard shadow MLPs for 10K iterations with a batch size of 256 in this stage, which takes about 7 hours for a scene.
Phase 3: BRDF and soft shadow field. In the section of material estimation (Sec. 3.5), the BRDF MLPs include four fully connected layers, each with 512 hidden units and ReLU activation functions. Following these four layers, the BRDF feature network is divided into separate albedo and roughness layers with 512 hidden units and Sigmoid activation. Specifically, albedo layer outputs the albedo term $\hat{A}\in\mathbb{R}^{3}$ and roughness layer outputs the roughness term $\hat{R}\in\mathbb{R}^{1}$ at surface position $\hat{\textbf{x}}$. We use 64 sample directions (8 $\times$ 16) over hemisphere $\Omega^{+}$ at sampled surface points for the specular environment map, and then when we use the specular environment map as specular incoming lighting with specular BRDF term $f_{s}$ to compute the rendering equation. We use important sampling to sample the incoming light radiance by interpolating the specular environment map based on Nvdiffrast [23], the important sampling number is 128.
@@ 0.B.2 Implemetaions of Baselines
Appendix 0.C Details of Experiments
0.C.1 Additional Results on Synthetic Dataset
We present additional qualitative (see Fig. 6 and 9) and quantitative (see Table 5,  6,  7,  8,  9,  10, and  11 comparison results on the synthetic dataset including 6 scenes. Compared to other inverse rendering methods, our approach excels in estimating albedo and decomposing shadows. Crucially, it effectively separates shadows and ensures that they are not incorrectly included in the albedo, especially under indoor lighting conditions. Note that PhySG [53] and InvRender [57] directly use our trained SDF-based radiance field as their geometry backbone. Since the predicted roughness of PhySG is global, it cannot converge on our indoor scene dataset. Therefore, we ignore the comparison of the roughness MSE metric with PhySG.
0.C.2 Additional Results on Real-World Dataset
Our comparative analysis on the real-world dataset, constrained to view synthesis metrics due to the unavailability of real scene material ground truths, is presented in Table 20. Although our method is slightly inferior to IBL-NeRF in terms of PSNR, it significantly outperforms other methods regarding SSIM and LPIPS metrics. Fig. 7 provides additional qualitative results that highlight our exceptional decomposition ability. We effectively disentangle albedo and shadow components, achieving physical plausibility in our results.
0.C.3 Additional Results for Ablation Studies
We showcase the effectiveness of our material optimization on synthetic datasets in Table 12131415161718, and 19. As described in Sec. 4.4 in the main paper, Without shadow terms significantly impacts albedo estimation, leading to incorrect albedo outputs that erroneously incorporate baked shadows; Without the soft shadow term, residual shadows remain baked into the albedo, and due to the binary nature of hard shadows, they fail to accurately represent physically plausible shadow elements; Without the incorporation of an instance-level albedo regularizer, there would likely be a lack of illumination consistency in the albedo for individual instances, potentially leading to less accurate synthesis results. We show more qualitative results in Fig. 8.
@@ 0.C.3 Additional Results for Ablation Studies
Figure 9: Qualitative results on Relighting of all methods on the synthetic dataset.
@@
[…]
Dinning Office Avg. Dinning Office Avg. Dinning Office Avg.
IBL-NeRF [4] 26.234 22.3221 24.2781 0.7968 0.7013 0.7491 0.4285 0.6165 0.5225